OpenStudy (anonymous):

l-3nl-2=4 a.2 or - 2 b.3 or - 3 c.4 or - 4 d.6 or - 6

OpenStudy (ash2326):

@ashleylynn did you try this ?

OpenStudy (ash2326):

Your previous question was also like this!!Try it, if you get stuck I'll help you:)

OpenStudy (anonymous):

I-3nI-2 +2=4+2 I-3nI =6 or I-3nI= -6 3n= 6 or 3n = -6 3n/3= 6/3 or 3n/3 = -6/3 n= 2 or n= -2

OpenStudy (anonymous):

did that help ?

OpenStudy (anonymous):

yes thank you so much!